Operations Report – March 19th

Operations Report 19-MAR-2020

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: none!

Notes on non-nominal systems: nothing to report

Wednesday

No longer using the generator. Solar Power only.

8 am SOC was 63

11 am SOC was 100

7 pm SOC was 97

Thursday

8 am SOC was 66

11:30 SOC was 68

4 pm SOC was 100

Diesel Reading – 66%

Station Propane Reading – 69 %

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): full gallons

Water Meter: n/a units

Water (static tank) – 150 gallons

Static to Loft Pump used – yes

Water in GreenHab – 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: no

Hours: 152.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: no

Hours: 104.5

Beginning charge: 100%

Ending charge: 100 %

Currently charging: yes

Spirit rover: no

Hours: 134.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: no

Hours: 70.3

Beginning charge: 100%

Ending charge: 100 %

Currently charging: yes

Curiosity rover used: no

Hours: 139.6

Beginning charge: 100%

Ending charge: 100 %

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: No

Reason for use: n/a

Oil Added? no

ATV Fuel Used: None

Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report.

HabCar used and why, where? Went to town for supplies

CrewCar used and why, where? CrewCar is currently in Hanksville and will remain there for a while.

General notes and comments: Remote power system monitoring is now online.

Summary of internet: I disconnected the wifi in the Hab and have the internet directly plugged into the remote power system online monitoring system. This means I can keep an eye on the system remotely.

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: GreenHab has been cleaned and prepped for summer except for the floor.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: This morning the SOC was 66% when the observatory was not being used, so the power loss we are experiencing at night is not caused by the observatory, but rather me. Observatory operations will continue unrestricted.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: We did not have any mud from the small amount of rain we got last night. Also, happy spring!